OpenStudy (anonymous):

If p represents "x = 7" and q represents "x2 = 49," which symbolic statement is true? Choose one answer. a. p → q b. q → p c. All of the above. d. p ↔ q

OpenStudy (jhannybean):

\[\large p : x=7\]\[\large q: x^2=49\] the left arrow means if, then. the double arrows mean?

OpenStudy (jhannybean):

if and only if?

OpenStudy (mathstudent55):

If x = 7, can you conclude correctly that x^2 = 49?

OpenStudy (anonymous):

$$p:x=7\\q:x^2=49$$So if you know \(x=7\), surely you can infer \(x^2=49\)? :-) For (b), we can't do the reverse; knowing \(x^2=49\) only lets us infer \(|x|=7\). Hence (c) and (d) are also wrong by proxy.